Consider the motion of particle mass m and charge q in an electromagnetic field with
electric field vector is E and the magnetic field vector is B. The force acting on the particle is
given by the Lorentz equation F = qE + qv x B (assuming non-relativistic case, v<<c).

(a) If there is no electric field and the particle enters the magnetic field in a direction
perpendicular to the lines of magnetic flux, show that the trajectory is a circle with radius
1
Expert's answer
2019-10-21T13:35:00-0400

Show that the trajectory is a circle with radius

We need to show the trajectory is a circle and find the radius of the circle

Solution:


From the given data = "\vec E"


vector of Electric field = "\vec B"


Charge in an Electro magnetic field = q


Force = "\vec F"


It is given by the Lorentz equation , that is

"\vec F = q \vec E + q \space (\vec \nu \space X \vec B)"


(a). If no Electric field is there, means "\vec E = 0"



So, the Lorentz equation becomes

"\vec F = q \space (\vec \nu \space X \vec B)"


Given information,

the particle enters the magnetic field ("\vec B" ), in perpendicular direction to the Lines of Magnetic Flux ("\vec V" )


It means, "\vec B \space and \space \vec V" are perpendicular





"\vec F =q\space (\vec V X \vec B)"

"\vec F = q \space V \space B \space sin \space \theta"


"Here, \theta = 90"


"\vec F = q \space V \space B \space"



Force is perpendicular to both "\vec V \space and \space \vec B" and it is in the direction of centre.

Here the magnitude of force is same but directions are different.

These are towards the centre.

So this makes the Trajectory as a Circle.


a force that acts on a body moving in a circular path and is directed towards the centre around which the body is moving

is called a "centripetal Force".



Centripetal force =

"\frac {m\nu^2} {r}"


Here r is the radius of the circle.



"q V B = \frac {m\nu^2} {r}"

"r = \frac {m \nu} {qB}"

Answer: shown that the trajectory is a circle and radius of the circle =

"r = \frac {m \nu} {qB}"